Circular No. 8866

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 N1 = P/2000 P3 (McNAUGHT) S. Nakano, Sumoto, Japan, has identified previously unpublished ob-

servations on three nights in 2000 August and November from the NEAT and LONEOS surveys with comet P/2007 N1 (cf. IAUC 8855, 8860). Following the announcement of this identification on MPEC 2007-R04, M. Meyer (Limburg, Germany) and R. J. Bouma (Groningen, The Netherlands) each independently found additional unrecognized NEAT images of the comet obtained with the Haleakala 1.2-m reflector from 2000 Sept. 3 and Dec. 19 — Meyer noting the comet to be diffuse with a tail, and Bouma writing that a stacked image from the Dec. exposures revealing a 4′′ coma and a faint tail ∼ 6′′ long in p.a. 45^{o} ± 5^{o}; Meyer’s astrometry appears on MPEC 2007-R17.

2007 DT_{103}

L. A. M. Benner, S. J. Ostro, J. D. Giorgini, J. Van Brimmer, L. Juare, J. S. Jao, R. F. Jurgens, and M. A. Slade, Jet Propulsion Laboratory, California Institute of Technology (CIT); and M. W. Busch, CIT, report that Goldstone (8560-mHz, 3.5-cm) radar observations during July 26, 28, and Aug. 1 reveal that minor planet 2007 DT_{103} (cf. MPEC 2007-F20; MPO 121877) is a binary system. Preliminary estimates for the component diameters are about 0.3 km and > 80 m. The maximum orbital distance between the components is at least 0.45 km.

(160256) 2002 PD_{149}

K. S. Noll and S. D. Kern, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; D. C. Stephens, Brigham Young University; and H. F. Levison, Southwest Research Institute, report the detection of a binary companion to the transneptunian object (160256) 2002 PD_{149} (MPEC 2002-S49; MPO 121459). The observations were made during May 22.195–22.212 UT with the Planetary Camera of the Wide Field Planetary Camera 2 on the Hubble Space Telescope, using the F606W filter (wide V) with one 260-s exposure at each of three dithered positions on the detector. The two components were separated by an angular distance of 0′′.74 ± 0′′.01 and differ in brightness by 0.4 mag. The fainter component lies at a position angle of 62.0 ± 0.8 deg from the primary. The projected separation of the objects in the sky plane is 24400 ± 300 km.

2007 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green